# Realizable Circuit Complexity: Embedding Computation in Space–Time

### v2

@@ 1 Introduction
2 Prior Work
Spatial and physical constraints on computation have received surprisingly little formal treatment in complexity theory. Nearly all foundational work in this direction traces back to a concise three-page note by Fisher [1]. Fisher modeled a $d$-dimensional parallel machine as a finite-density grid of processors communicating at bounded signal speed. He proved that any computation performed within time $t$ must lie inside a causal sphere of radius $r=\Theta(t)$ and, consequently, that the achievable execution time for a computation with $I$ inputs, $K$ outputs, and total processor–time $T$ satisfies
$t=\Omega\big(\max\{\,I^{1/d},\,K^{1/d},\,T^{1/(d+1)}\,\}\big),$ (1)
Subsequent efforts, most notably the Spatial Machine framework of Leighton, Rose, and Saxe [2] and related analyses [3, 4], extended Fisher’s geometric intuition. These models formalize computation as a grid of finite-state processors, each capable of communicating only with local neighbors. They confirm that the physical embedding of computation in finite-dimensional space imposes fundamental communication and latency limits beyond those captured by traditional circuit complexity.
Limitations of Fisher’s formulation.
$t=\Omega(n^{1/(d-1)}),$ (2)
which is a strictly tighter lower bound than (1) for $d>2$.
Beyond geometry: flux and thermodynamics.
  1. 1.
  2. 2.
  3. 3.
Toward a physically grounded complexity framework.
@@ 3.3 Impracticality
    Polynomial gate count as a feasibility proxy. In $\mathbf{P/poly}$ and its subclasses, feasibility is defined by the existence of circuits with polynomially many gates. However, the polynomial degree can mask exponential resource blow-up in area, wiring, or latency. VLSI area–time tradeoffs [6] show that global operations such as sorting or FFT require $A\cdot T^{2}=\Omega(N^{2})$ resources, even when gate count is modest. Geometry and interconnects dominate cost, making gate count alone a poor indicator of realizability.
@@
    Unbounded fan-in. The classes $\mathbf{AC}$ and $\mathbf{TC}$ permit unbounded fan-in gates, effectively assuming that arbitrarily many signals can be aggregated within a single layer. Foundational VLSI results [5] and empirical wiring laws [7, 8, 11] demonstrate that I/O pins and interconnects scale sub-linearly with logic, imposing strict upper bounds on fan-in and fan-out. Thus, unbounded fan-in is not physically realizable.
@@
    Neglect of energy and data movement. Data movement dominates energy consumption in modern computing [9, 10]. Because $\mathbf{NC/AC/TC}$ ignore interconnects and memory-transfer energy, they overestimate the scalability of shallow-depth circuits that depend on global aggregation or broadcasting.
@@
4 Approach
A physically meaningful theory of circuits must account for the constraints that govern their realization in space and time. Thus, our primary goal is to closely align the theoretical treatment of circuits with the physical.
4.1 Physical constraints of computation
We propose three fundamental bounds on circuit proportions, motivated by principles from physics and information theory. In Section 5, we provide a formal derivation from a physical model of computation. (For a consideration of application to quantum systems, please refer to Appendix B.)
@@ 5 Formalism for Physical Constraints
To formalize these constraints, we model physical computation as local Hamiltonian systems evolving inside causal regions. This ensures our abstraction assumes only unavoidable physical principles: finite signal speed, locality, and conservation. In particular, the $\mathcal{O}(t^{d})$ volume bound follows from comparison geometry (Bishop–Gromov [12]) under nonnegative Ricci curvature, while the $O(t^{d-1})$ flux bound follows from Liouville’s theorem on Hamiltonian flows.
@@ 5.1 Volume bound
Notice in Definition 5.2 that geodesic distance allows for curved spatial geometry. When space is flat, $d_{g}(\cdot,\cdot)$ is the usual Euclidean metric. When positive Ricci curvature is present, it follows from comparison geometry (Bishop–Gromov [12]) that the volume within a boundary is strictly smaller than in the Euclidean case.
@@ 5.3 Density bound
6 The $\mathbf{RC}_{d}$ Framework
@@ 6.1 Realizable Circuits
6.2 Case study: Attention and transformer scaling.
We demonstrate the implications of $\mathbf{RC}_{d}$ through a detailed analysis of attention mechanisms (Appendix A). By mapping multi-head attention to the flux-bounded dynamics of a realizable circuit, we show that all attention variants—softmax, average-hard, kernelized, or sparse—obey a universal spatiotemporal scaling law:
$T\;=\;\Omega\big(I^{\star}(\varepsilon;n)^{1/d}\big),$
where $I^{\star}(\varepsilon;n)$ is the cross-cut information demand of the task. This bound follows directly from the flux conservation theorem ($\dot{I}\sim t^{d-1}$) integrated over causal time, and explains why increasing head count $H$ improves only constant factors ($H^{-1/d}$) but cannot alter the exponent. The $\mathbf{RC}_{d}$ framework therefore refines classical $\mathbf{TC^{0}}$ analyses of transformers by revealing the physical origin of the observed sub-linear scaling: realizable circuits are inherently limited by geometric and thermodynamic flux constraints, not merely symbolic depth or fan-in.
6.3 Realizable Circuits in the limit
@@ 7.3 In the circuit complexity hierarchy
By condition (S) of Definition 6.1, every function in $\mathbf{RC}_{d}(t(n))$ is computable by a circuit of size $\mathcal{O}(t(n)^{d})$ and bounded fan-in. Thus, it lies in $\mathbf{SIZE}(\mathcal{O}(t(n)^{d}))$. If $t(n)$ is polynomial, the circuit size is also polynomial, placing the family within $\mathbf{P/poly}$. The inclusion of $\mathbf{NC}$ follows from Theorem 6.5. ∎
@@
8 Analysis
8.1 Relation to Landauer’s principle
The dissipation of heat is an essential physical limit on computation but does not alter the asymptotic bounds already imposed by geometry in $\mathbf{RC}_{d}$. It instead reinforces them by constraining how fast logically irreversible operations can occur within a finite causal region. Landauer’s principle states that information erasure has a thermodynamic cost, ensuring that entropy balance is preserved even when logical information is discarded.
Definition 8.1 (Landauer’s principle).
$E_{\min}=k_{B}T\ln 2$
of heat energy into the surrounding environment.
$Q\ \geq\ k_{B}T(\ln 2)\,\Delta I.$
Theorem 8.2 (Thermodynamic constraint on realizable circuits).
$k_{B}T(\ln 2)\,E(t)\ \leq\ Q(t)\ \leq\ \eta\,t^{d-1},$
for some device-dependent constant $\eta$. Consequently,
[…]
Proof.
Remark (Connection to reversible computation).
Landauer’s principle implies that only logically irreversible operations incur an unavoidable thermodynamic cost. Bennett [13] showed that any classical computation can be made logically reversible by preserving all intermediate information until the output is copied, after which the computation can be uncomputed to erase ancillary states without dissipating heat. In practice, however, reversibility transfers the cost from heat to space: each bit preserved increases the causal size of the circuit.
8.2 On extension to recurrent circuits
Definition 8.3 (Recurrent realization).
$F_{C}:\{0,1\}^{n}\to\{0,1\}^{n},\qquad w(t+1)=F_{C}(w(t)).$
Theorem 8.4 (Temporal completeness at $d=1$).
Proof.
Corollary 8.5 (Dimensional completeness of $\mathbf{RRC}_{d}$).
Remark (Interpretation).
@@ A Relation to Transformers
Connection to circuit complexity. Although many computational domains can be analyzed through realizability constraints, we emphasize their implications for machine-learning architectures. Historically, the circuit class $\mathbf{TC}$ was introduced by Parberry and Schnitger (1988) [14] to capture the threshold-gate expressivity of early neural networks. Our formulation extends this lineage by defining $\mathbf{RC}_{d}$ as a physically realizable subclass of circuit families, incorporating spatial locality, finite signal velocity, and thermodynamic limits. Under this model, modern transformer layers—which consist of parallel attention heads, feed-forward sublayers, and residual routing—correspond to a bounded-flux computation embedded in three-dimensional space (d=3d{=}3).
@@ A.1 Prior results on transformers and circuit complexity
Many non-recurrent neural networks—including transformers—can be modeled as highly parallel threshold circuits of constant depth. Merrill and Sabharwal (2023) [15] prove that average-hard attention transformers ($\mathbf{AHAT}$) and softmax attention transformers ($\mathbf{SMAT}$) lie in $\mathbf{TC^{0}}$, with $\mathbf{SMAT}$ satisfying a $\mathbf{DLOGTIME}$-uniformity condition [16]. Strobl (2023) [17] further tightens $\mathbf{AHAT}$ by imposing an $\mathbf{L}$-uniformity constraint, while Chiang (2025) [18] establishes $\mathbf{AHAT}=\mathbf{SMAT}$ by achieving the same $\mathbf{DLOGTIME}$-uniformity bound. For completeness, we also include the earlier result of Hahn (2020) [19], which upper-bounds the class of hard attention transformers ($\mathbf{HAT}$) by $\mathbf{AC^{0}}$.
@@
Hard Attention ($\mathbf{HAT}$) $\mathbf{AC^{0}}$ [19]
Softmax Attention ($\mathbf{SMAT}$) $\mathbf{AC^{0}}$ [15] $\mathbf{DLOGTIME}$-uniform $\mathbf{TC^{0}}$ [16]
Average-Hard Attention ($\mathbf{AHAT}$) $\mathbf{AC^{0}}$ [15] $\mathbf{DLOGTIME}$-uniform $\mathbf{TC^{0}}$ [18]
@@ A.5 Attention under joint throughput and Landauer limits in $\mathbf{RC}_{d}$
B Applicability to Quantum Computation
@@
C Position in the History of Computational Theory

### v3

@@ 2.3 Impracticality
    Polynomial gate count as a feasibility proxy. In $\mathbf{P/poly}$ and its subclasses, feasibility is defined by the existence of circuits with polynomially many gates. However, the polynomial degree can mask exponential resource blow-up in area, wiring, or latency. VLSI area–time tradeoffs [Thompson1979ATVLSI] show that global operations such as sorting or FFT require $A\cdot T^{2}=\Omega(N^{2})$ resources, even when gate count is modest. Geometry and interconnects dominate cost, making gate count alone a poor indicator of realizability.
@@
    Unbounded fan-in. The classes $\mathbf{AC}$ and $\mathbf{TC}$ permit unbounded fan-in gates, effectively assuming that arbitrarily many signals can be aggregated within a single layer. Foundational VLSI results [MeadConway1980VLSI] and empirical wiring laws [LandmanRusso1971Rent, Lanzerotti2005RentIBMJRD, Stroobandt2010RentRuleSurvey] demonstrate that I/O pins and interconnects scale sub-linearly with logic, imposing strict upper bounds on fan-in and fan-out. Thus, unbounded fan-in is not physically realizable.
@@
    Neglect of energy and data movement. Data movement dominates energy consumption in modern computing [Horowitz2014EnergyProblem, SzeNASEM2024DataMovement]. Because $\mathbf{NC/AC/TC}$ ignore interconnects and memory-transfer energy, they overestimate the scalability of shallow-depth circuits that depend on global aggregation or broadcasting.
@@
2.4 Prior work
Spatial and physical constraints on computation have received surprisingly little formal treatment in complexity theory. Nearly all foundational work in this direction traces back to a concise three-page note by Fisher [Fisher1988]. Fisher modeled a $d$-dimensional parallel machine as a finite-density grid of processors communicating at bounded signal speed. He proved that any computation performed within time $t$ must lie inside a causal sphere of radius $r=\Theta(t)$ and, consequently, that the achievable execution time for a computation with $I$ inputs, $K$ outputs, and total processor–time $T$ satisfies
$t=\Omega\big(\max\{\,I^{1/d},\,K^{1/d},\,T^{1/(d+1)}\,\}\big),$ (1)
Subsequent efforts, most notably the Spatial Machine framework of Leighton, Rose, and Saxe [spatial_machines] and related analyses [JEN1989331, WORLEY1992361], extended Fisher’s geometric intuition. These models formalize computation as a grid of finite-state processors, each capable of communicating only with local neighbors. They confirm that the physical embedding of computation in finite-dimensional space imposes fundamental communication and latency limits beyond those captured by traditional circuit complexity.
Limitations of Fisher’s formulation.
$t=\Omega(n^{1/(d-1)}),$ (2)
which is a strictly tighter lower bound than (1) for $d>2$.
Beyond geometry: flux and thermodynamics.
  1. 1.
  2. 2.
  3. 3.
Toward a physically grounded complexity framework.
3 Physical Constraints
We propose three fundamental bounds on circuit proportions, motivated by principles from physics and information theory. In Section 4, we provide a formal derivation from a physical model of computation. (For a consideration of application to quantum systems, please refer to Appendix B.1.)
@@ 4 Formalism for Physical Constraints
To formalize these constraints, we model physical computation as local Hamiltonian systems evolving inside causal regions. This ensures our abstraction assumes only unavoidable physical principles: finite signal speed, locality, and conservation. In particular, the $\mathcal{O}(t^{d})$ volume bound follows from comparison geometry (Bishop–Gromov [Bishop1964GeometryOM]) under nonnegative Ricci curvature, while the $O(t^{d-1})$ flux bound follows from Liouville’s theorem on Hamiltonian flows.
@@ 4.1 Volume bound
Notice in Definition 4.2 that geodesic distance allows for curved spatial geometry. When space is flat, $d_{g}(\cdot,\cdot)$ is the usual Euclidean metric. When positive Ricci curvature is present, it follows from comparison geometry (Bishop–Gromov [Bishop1964GeometryOM]) that the volume within a boundary is strictly smaller than in the Euclidean case.
@@ 4.3 Density bound
4.4 Thermodynamics connection
The dissipation of heat is an essential physical limit on computation but does not alter the asymptotic bounds already imposed by geometry in $\mathbf{RC}_{d}$. It instead reinforces them by constraining how fast logically irreversible operations can occur within a finite causal region.
4.4.1 Relation to Landauer’s principle
Landauer’s principle states that information erasure has a thermodynamic cost, ensuring that entropy balance is preserved even when logical information is discarded.
Definition 4.14 (Landauer’s principle).
$E_{\min}=k_{B}T\ln 2$
of heat energy into the surrounding environment.
$Q\ \geq\ k_{B}T(\ln 2)\,\Delta I.$
Theorem 4.15 (Thermodynamic constraint on realizable circuits).
$k_{B}T(\ln 2)\,E(t)\ \leq\ Q(t)\ \leq\ \eta\,t^{d-1},$
for some device-dependent constant $\eta$. Consequently,
[…]
Proof.
Remark (Connection to reversible computation).
Landauer’s principle implies that only logically irreversible operations incur an unavoidable thermodynamic cost. Bennett [Bennett1973LogicalRC] showed that any classical computation can be made logically reversible by preserving all intermediate information until the output is copied, after which the computation can be uncomputed to erase ancillary states without dissipating heat. In practice, however, reversibility transfers the cost from heat to space: each bit preserved increases the causal size of the circuit.
4.4.2 Relation to thermodynamic circuit frameworks
A line of work in non-equilibrium statistical physics quantifies the energetic and entropic costs of computation realized by circuits. Wolpert and Kolchinsky analyze how gate topology and physical coupling determine the thermodynamic costs of straight-line circuits [WolpertKolchinsky2020ThermoCircuits], while Kolchinsky and Wolpert bound entropy production and extractable work under protocol constraints for general state transformations [KolchinskyWolpert2021PRX]. More recently, Yadav, Yousef, and Wolpert develop the mismatch cost (MMC) formalism to lower-bound the minimal entropy production per run of a circuit, relate MMC to circuit size, and compare families implementing the same Boolean function [YadavYousefWolpert2025MMC]. These frameworks rigorously connect logical structure to thermodynamic cost, largely independent of spatial geometry.
By contrast, $\mathbf{RC}_{d}$ is a geometric–causal refinement of circuit complexity: it imposes explicit embedding in $\mathbb{R}^{d}$, a causal light-cone, and a cut-set (surface) bound on information throughput.
Assumptions.
Thermodynamic circuit models assume stochastic (often Markovian) dynamics and analyze entropy production for given gate couplings/topologies, without fixing an ambient spatial dimension or causal cone [WolpertKolchinsky2020ThermoCircuits, KolchinskyWolpert2021PRX]. The MMC framework adds distributional mismatch between gate priors and actual inputs to quantify minimal EP per evaluation [YadavYousefWolpert2025MMC]. In contrast, $\mathbf{RC}_{d}$ posits a finite-density embedding in $d$ dimensions with maximum signal speed and derives two geometric constraints: volume $|C(t)|=\mathcal{O}(t^{d})$ and boundary throughput $w(t)=\mathcal{O}(t^{d-1})$. These imply a cut-set capacity law after integrating flux over time, independent of any particular stochastic micro-model.
Primary objects of study.
The thermodynamic works bound entropy production (EP), work, and minimal thermodynamic cost as a function of circuit topology and protocol; they treat energy/EP as the key resources [KolchinskyWolpert2021PRX, WolpertKolchinsky2020ThermoCircuits]. $\mathbf{RC}_{d}$ instead treats space–time throughput as the primitive: it constrains the number of distinguishable bits that can cross any geometric cut per unit time by a surface law, and only then invokes Landauer to convert lost accessible information into a compatible heat bound.
Results comparison.
  • Thermodynamic frameworks derive EP/work lower bounds that depend on circuit topology, priors, and protocol constraints, including cases where MMC scales with circuit size or diverges for mismatched priors [KolchinskyWolpert2021PRX, YadavYousefWolpert2025MMC]. These do not by themselves yield a surface-law limit on information throughput or a $t^{d-1}$ bound tied to spatial dimension.
  • $\mathbf{RC}_{d}$ proves geometry-driven flux bounds: instantaneous cross-boundary flow is $\mathcal{O}(t^{d-1})$ and total cross-cut information over time $T$ is $\Theta(T^{d})$. This yields universal scaling limits (e.g., $T\geq\Omega(n^{1/d})$) for any realizable computation (classical, quantum, encrypted) and enables cut-set lower bounds for attention and FHE that are independent of particular device physics.
Complementarity.
The frameworks are complementary: thermodynamic formalisms quantify how much entropy must be produced by a given (possibly geometry-agnostic) circuit and protocol; $\mathbf{RC}_{d}$ quantifies how fast accessible information and heat can cross a causal boundary in $d$-dimensional space. Combining them yields the joint feasibility region used in our Landauer–throughput theorems: the EP required by MMC (or protocol constraints) must be exported no faster than the $\mathbf{RC}_{d}$ flux allows, enforcing a time lower bound $T$ with exponent $1/d$.
Scope and novelty.
Where prior thermodynamic analyses focus on EP/work optima for given circuit topologies and protocols [WolpertKolchinsky2020ThermoCircuits, KolchinskyWolpert2021PRX, YadavYousefWolpert2025MMC], $\mathbf{RC}_{d}$ introduces a dimension-indexed hierarchy of realizable circuits with universal geometric bounds that persist across computational substrates. This enables new, geometry-tight limits for highly parallel mechanisms (e.g., multi-head attention) and privacy-preserving computation (FHE), which are not addressed by MMC or protocol-constraint formalisms alone.
4.5 Case study: Attention and transformer scaling.
We demonstrate the implications of our framework through a detailed analysis of attention mechanisms (Appendix A). By mapping multi-head attention to the flux-bounded dynamics of a realizable circuit, we show that all attention variants—softmax, average-hard, kernelized, or sparse—obey a universal spatiotemporal scaling law:
$T\;=\;\Omega\big(I^{\star}(\varepsilon;n)^{1/d}\big),$
where $I^{\star}(\varepsilon;n)$ is the cross-cut information demand of the task. This bound follows directly from the flux conservation theorem ($\dot{I}\sim t^{d-1}$) integrated over causal time, and explains why increasing head count $H$ improves only constant factors ($H^{-1/d}$) but cannot alter the exponent. The Realizable Circuits framework therefore refines classical $\mathbf{TC^{0}}$ analyses of transformers by revealing the physical origin of the observed sub-linear scaling: realizable circuits are inherently limited by geometric and thermodynamic flux constraints, not merely symbolic depth or fan-in.
5 The $\mathbf{RC}_{d}$ Family
@@ 5.1 Realizable Circuits
5.1.1 Extensions
It is important to note that we choose our definition of $\mathbf{RC}_{d}$ to match the established treatment of circuits. In particular, $\mathbf{RC}_{d}$ is defined in terms of classical information and also prohibits any recurrent connections. However, our physical framework allows us to model methods of computation that extend beyond this limited scope. For a discussion on extension of $\mathbf{RC}_{d}$ to quantum and recurrent circuits, please refer to Appendix B.
5.1.2 Realizable Circuits in the limit
@@ 6.3 In the circuit complexity hierarchy
By condition (S) of Definition 5.1, every function in $\mathbf{RC}_{d}(t(n))$ is computable by a circuit of size $\mathcal{O}(t(n)^{d})$ and bounded fan-in. Thus, it lies in $\mathbf{SIZE}(\mathcal{O}(t(n)^{d}))$. If $t(n)$ is polynomial, the circuit size is also polynomial, placing the family within $\mathbf{P/poly}$. The inclusion of $\mathbf{NC}$ follows from Theorem $5.$ ∎
@@ A Relation to Transformers
Connection to circuit complexity. Although many computational domains can be analyzed through realizability constraints, we emphasize their implications for machine-learning architectures. Historically, the circuit class $\mathbf{TC}$ was introduced by Parberry and Schnitger (1988) [PARBERRY1988278] to capture the threshold-gate expressivity of early neural networks. Our formulation extends this lineage by defining $\mathbf{RC}_{d}$ as a physically realizable subclass of circuit families, incorporating spatial locality, finite signal velocity, and thermodynamic limits. Under this model, modern transformer layers—which consist of parallel attention heads, feed-forward sublayers, and residual routing—correspond to a bounded-flux computation embedded in three-dimensional space (d=3d{=}3).
@@ A.1 Prior results on transformers and circuit complexity
Many non-recurrent neural networks—including transformers—can be modeled as highly parallel threshold circuits of constant depth. Merrill and Sabharwal (2023) [merrill2022saturated] prove that average-hard attention transformers ($\mathbf{AHAT}$) and softmax attention transformers ($\mathbf{SMAT}$) lie in $\mathbf{TC^{0}}$, with $\mathbf{SMAT}$ satisfying a $\mathbf{DLOGTIME}$-uniformity condition [merrill2023unisoft]. Strobl (2023) [strobl2023average] further tightens $\mathbf{AHAT}$ by imposing an $\mathbf{L}$-uniformity constraint, while Chiang (2025) [chiang2024transformers] establishes $\mathbf{AHAT}=\mathbf{SMAT}$ by achieving the same $\mathbf{DLOGTIME}$-uniformity bound. For completeness, we also include the earlier result of Hahn (2020) [hahn-2020-theoretical], which upper-bounds the class of hard attention transformers ($\mathbf{HAT}$) by $\mathbf{AC^{0}}$.
@@
Hard Attention ($\mathbf{HAT}$) $\mathbf{AC^{0}}$ [hahn-2020-theoretical]
Softmax Attention ($\mathbf{SMAT}$) $\mathbf{AC^{0}}$ [merrill2022saturated] $\mathbf{DLOGTIME}$-uniform $\mathbf{TC^{0}}$ [merrill2023unisoft]
Average-Hard Attention ($\mathbf{AHAT}$) $\mathbf{AC^{0}}$ [merrill2022saturated] $\mathbf{DLOGTIME}$-uniform $\mathbf{TC^{0}}$ [chiang2024transformers]
@@ A.5 Attention under joint throughput and Landauer limits in $\mathbf{RC}_{d}$
B Future Extensions
The framework of $\mathbf{RC}_{d}$ is not confined to Boolean or quantum circuits; it provides a general account of physically realizable information processing. Because its axioms arise from geometry (finite density), causality (bounded propagation), and thermodynamics (finite entropy flux), the same bounds apply to any substrate that computes through physical interaction and state evolution.
B.1 On extension to quantum circuits
@@
B.2 On extension to recurrent circuits
Definition B.4 (Recurrent realization).
$F_{C}:\{0,1\}^{n}\to\{0,1\}^{n},\qquad w(t+1)=F_{C}(w(t)).$
Theorem B.5 (Temporal completeness at $d=1$).
Proof.
Corollary B.6 (Dimensional completeness of $\mathbf{RRC}_{d}$).
Remark (Interpretation).
B.3 Other domains
Finally, we highlight how these constraints manifest across diverse computational domains.
Neuromorphic, embodied, and mortal computation.
Neural and embodied systems compute through continuous, spatially extended dynamics constrained by finite conduction speed and local connectivity. Within $\mathbf{RC}_{3}$, the flux limit $w(t)=\mathcal{O}(t^{2})$ imposes a ceiling on global synchronization and long-range coordination, forcing large-scale inference to emerge hierarchically. This constraint renders cortical and neuromorphic architectures—notably predictive-coding [rao1999predictive, salvatori2025survey] and columnar models [Horton2005Cortical, Buzsaki2006Rhythms, Sporns2016ModularBN, Eliasmith2013Nengo]—realizable in the physical sense: they operate near the entropy–flux bound of $\mathbf{RC}_{3}$. From this perspective, embodied intelligence [gupta2021embodied] represents computation that is spatially embedded within causal geometry, while mortal computation [ororbia2023mortal] denotes the same process under finite energy and entropy budgets. Both are natural instantiations of realizable computation, grounded in the same physical limits of signal propagation and entropy flux.
Quantum and reversible computation.
Quantum systems obey the same geometric constraints while preserving information reversibly. The inclusion $\mathbf{RC}_{d}\subseteq\mathbf{QRC}_{d}$ follows from the existence of unitary embeddings for any classical process, but the converse fails because irreversible logic cannot in general be uncomputed. Even quantum systems are flux-bounded by the Lieb–Robinson limit [LiebRobinson1972, Nachtergaele2006LiebRobinson], which restricts information propagation to a linear light-cone. Thus $\mathbf{QRC}_{d}$ refines $\mathbf{BQP}$ by adding explicit spatial and flux constraints, clarifying why even quantum hardware cannot surpass the $t^{d-1}$ throughput limit [Bravyi2006LiebRobinsonQuantum, bekenstein1990quantum].
VLSI and spatial hardware design.
For planar devices ($d=2$), the $\mathbf{RC}_{2}$ bound recovers the area–time tradeoff $AT^{2}=\Omega(N^{2})$ of Thompson’s VLSI theory [Thompson1979ATVLSI] and Rent’s empirical rule on interconnect scaling [LandmanRusso1971Rent, Stroobandt2010RentRuleSurvey]. Here $A=\mathcal{O}(t^{2})$ corresponds to physical area, while the boundary throughput $w(t)=\mathcal{O}(t)$ constrains wiring density. Hence $\mathbf{RC}_{d}$ subsumes classical VLSI theory as its two-dimensional limit and naturally extends it to higher-dimensional substrates [MeadConway1980VLSI].
Distributed, edge, and communication systems.
In distributed computing or sensor networks embedded in $\mathbb{R}^{d}$, information exchange between nodes is limited by the same causal and flux constraints. The $\mathcal{O}(t^{d-1})$ bound on cross-boundary transfer defines a geometric analogue of Shannon capacity [shannon1948mathematical], setting a hard limit on aggregate throughput. This result parallels Gupta and Kumar’s spatial capacity law for wireless networks, where total achievable rate scales as $O(\sqrt{n})$ in planar space [gupta2002capacity]. Thus $\mathbf{RC}_{d}$ unifies circuit and network theories through a common geometric formulation of communication.
Energy-aware and thermodynamic machine learning.
When learning dynamics are interpreted as physical state updates, Liouville’s theorem and Gibbs entropy conservation imply that the energy–delay product of any model is lower-bounded by its entropy flux. Consequently, the description-length or free-energy minimization objectives in modern energy-based and predictive-coding models [friston2010free, scellier2017equilibrium, alemi2016deep, tishby2015deep, salvatori2025survey] are realizable precisely when their updates satisfy the $\mathbf{RC}_{d}$ bounds. This connects the Landauer limit [landauer1961irreversibility, bennett1982thermodynamics], information geometry, and computational generalization under a single physical principle.
Unified scaling law.
$t=\Omega\big(n^{1/(d-1)}\big),\qquad w(t)=\mathcal{O}(t^{d-1}),$
C Final Remarks